# Breaking through the learning plateaus of in-context learning in Transformer

### v2

@@ 4.3 Disfunction of weights component
Confusing pattern of weights component  We executed the task under two conditions: $D_{\text{fix}}\Rightarrow D_{\text{fix}}$ and $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$, with the outcomes presented in Fig. 4A. As expected, in the simpler scenario of $D_{\text{fix}}\Rightarrow D_{\text{fix}}$, both the context and weights components exhibit improvement throughout the learning process, leading to enhanced in-context learning performance. However, in the more challenging setting of $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$, the weights component deteriorates over time, with its score remaining below the initial value for the entire duration of the learning process. This differs from the context component, which improves with the rise in in-context learning performance. We refer to the situation where the weights component score falls below the starting value as a disfunction of the weights component. This outcome is intriguing because training has no effect in improving the weights component. To gain better insight into this phenomenon, we carried out additional experiments across varying levels of task complexity. The model was trained for 50 epochs on these tasks, and we monitored the weights component score post-training. We found that the weights component score gradually declines as the entropy increasing, eventually stabilizing at around 0.8, as shown in Fig. 4B.
@@
Weight component degradation is linked to duration learning plateaus.  Our primary concern is the duration of learning plateaus, and we seek to comprehend its connection to the in-weights component. To investigate this, we graphed the relationship between learning plateau length and weights component score after 50 epochs, as shown in Fig. 4C. Our analysis reveals an approximate linear correlation between the weights component score and the learning plateau duration. Short learning plateaus occur when there is significant improvement in the in-weights component score. Conversely, long learning plateaus arise when the weights component is disfunctioning or on the edge of disfunction, that is, when the weights component score is at or below the initial value.
Figure 4: Weights component and learning plateaus. A. The weights component score is increasing under $D_{\text{fix}}\Rightarrow D_{\text{fix}}$, while the weights component score is descreasing under the $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$ setting. Note that the “Weights” and “Context” in the figure is short for weights comp. score and context comp. score respectively. B. The weights component score after 50 epoch training is decrease when increase the complexity of task. The dash green line indicate the weights component score at initialization point. C. The weights component score at 50 epoch is negative correlates with the length of learning plateaus.
@@
Why the weights component is related to learning plateaus.  We are attempting to comprehend this phenomenon, yet the theoretically analyzing the learning process of a multilayer transformer on intricate data proves to be a formidable challenge. Recent theoretical studies [42, 13, 21] focus on the learning dynamics of Transformers with one or two layers using simple datasets. Given these limitations, we propose a more attainable, albeit weaker, construction analysis in Appendix E. Our approach is grounded in the notion that if a model with a good weights component can enhance its in-context learning capabilities with just a few additional parameters, then the weights component must be pivotal for achieving in-context learning prowess. To test this idea, we hypothesize that the in-weights component has been perfectly learned within a specific layer of Transformer. Our findings reveal that by adding at most three extra Transformer layers specifically tailored for processing contextual information, the model demonstrates significant in-context learning performance. These outcomes suggest that in-context learning abilities are more readily achieved when the weights component is well-optimized.
@@ 5 Breaking Through Learning Plateaus
The previous section examined the mechanisms behind learning plateaus and identified their connection to the weights component. In this section, our goal is to investigate whether we can shorten the learning plateaus or improve the performance increasing for each transition process without scaling models. We have chosen the $D_{\text{rnd}}$ configure as the configuration of the test set by default, as it has demonstrated pronounced learning plateau behavior, and this particular scenario is known to be more effective in evaluating in-context learning capabilities, as discussed by Wei et al. [48], Min et al. [32].
@@
Figure 5: Three methods proposed to assist in overcoming learning plateaus. A: Effective of the warm up method. Top: Employing $D_{\text{fix}}$ as a warm-up for the Transformer significantly mitigates learning plateaus. The dashed line indicates the transition point from $D_{\text{fix}}$ to $D_{\text{rnd}}$. Bottom: We execute the transition from $D_{\text{fix}\to\text{rnd}}\Rightarrow D_{\text{rnd}}$ at various switching points. The curve labeled "2" signifies the switch from $D_{\text{fix}}$ to $D_{\text{rnd}}$ at epoch 2. The curve labeled "0" serves as the baseline, that is, $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$. The dashed lines highlight the respective switching points. B: Combining $D_{\text{fix}}$ and $D_{\text{rnd}}$. Top: Mixed training substantially improves the weights component score during the learning process and eliminates learning plateaus. Bottom: Boosting the weights component can promote the development of in-context learning capabilities in smaller models. The dashed line depicts the task configuration $D_{\text{fix}\land\text{rnd}}\Rightarrow D_{\text{rnd}}$, while the solid line represents the $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$ setting. C: Extra Loss. Top: Incorporating an weights loss can significantly enhance learning, whereas adding context loss does not have a noticeable impact. The baseline is $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$. Bottom: With the weights loss, the Transformer can attain a commendable weights component score after 50 epochs of training. The green dashed line indicates the weights component score at the outset.
5.1 Weights warm up method
@@
Previous experiments demonstrate that the model better learns the weight component more effectively in the $D_{\text{fix}}$ setting. A straight forward intuition is that whether we can use $D_{\text{fix}}$ to improve the weights component in $D_{\text{rnd}}$ test set. We give a further analysis about the relation between $D_{\text{rnd}}$ and $D_{\text{fix}}$:
@@
For $D_{\text{fix}}$: Recall that we have ${\mathbb{P}}(y_{p}|{\mathbf{x}}_{p},s_{c})\sim{\mathbb{P}}(v_{p}|{\mathbf{x}}_{p}){\mathbb{P}}(e_{h}|s_{c},m){\mathbb{P}}(m|s_{c})$. Under $D_{\text{fix}}$ setting, we only has one mapping function $m_{0}$. Consequently, our model can readily learn that ${\mathbb{P}}(m_{0}|s_{c})=1$. This simplification allows the model to concentrate on mastering ${\mathbb{P}}(v_{\mathbf{x}}|{\mathbf{x}})$. Therefore, it is anticipated that the model will develop a more refined weights component under this setting. Knowledge transfering between $D_{\text{fix}}$ and $D_{\text{rnd}}$ settings: The knowledge of ${\mathbb{P}}(v_{\mathbf{x}}|{\mathbf{x}})$ is shared between these two settings. The reason is that ${\mathbb{P}}(v_{\mathbf{x}}|{\mathbf{x}})$ is unrelated to the mapping function $m$ and the context $s_{c}$.
@@
Based on this analysis, we propose the following data configuration to improve the weights component on $D_{\text{fix}}$ before training on $D_{\text{rnd}}$. This setting is denoted as $D_{\text{fix}\to\text{rnd}}$, which means that we initially train the model on $D_{\text{fix}}$ for a specific epoch (weights warm up), and then, we train the model on $D_{\text{rnd}}$. We have the following discovery based on the experiments on $D_{\text{fix}\to\text{rnd}}\Rightarrow D_{\text{rnd}}$.
@@
Weights warm up help mitigate the learning plateaus. Fromv Fig. 5A Top, we observe a notable enhancement in the weights component during the initial “warm up” phase. Subsequent to this phase, there is a swift improvement in the accuracy of the in-context learning. Nonetheless, a subsequent decline in the weights component score post-warm up is apparent. This further verify that the Transformer can not learn a good weights component under $D_{\text{rnd}}$ settings. Further investigation is conducted by training the Transformer with various “warm up” durations under the $D_{\text{fix}}$ setting. As shown in Figure5A Bottom, training the Transformer for a brief number of epochs effectively eliminates the initial learning plateaus. Furthermore, this “warm up” phase contributes to a reduction in the duration of the subsequent learning plateaus. The application of the “warm up” technique yields significantly better results compared to those without it, particularly after a training duration of 50 epochs.
5.2 Mix training method
@@
In this section, we try to improve the weights component during the whole training process by designing the $D_{\text{fix}\land\text{rnd}}\Rightarrow D_{\text{rnd}}$ setting. In $D_{\text{fix}\land\text{rnd}}$ setting, half training data comes from $D_{\text{fix}}$ setting and half training data comes from $D_{\text{rnd}}$ setting. For fair comparison, the total number of training data is the same as the training set that using $D_{\text{rnd}}$ or $D_{\text{fix}}$ setting only.
@@
The mix training method significant boost the learning process.  As anticipated, both the weights component and the task accuracy exhibit improvements throughout the learning process, as depicted in Fig.5B Top. And the results do not show a pronounced learning plateau. This suggests that concurrently enhancing weights component is beneficial. In Fig.5B Bottom, we compare models of various sizes trained on the $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$ setting (represented by the solid line) against those trained on the $D_{\text{rnd}\land\text{fix}}\Rightarrow D_{\text{rnd}}$ setting (depicted with a dashed line). It is observed that a model of size 6.69M trained on $D_{\text{fix}\land\text{rnd}}$, after 20 epochs of training, can attain an accuracy comparable to a model of size 13.88M trained on $D_{\text{rnd}}$ that has undergone 50 epochs of training. These findings suggest that overcoming learning plateaus can lead to a reduction in the computational resources required for training.
@@ 5.3 Extra loss method
Figure 6: Experiments on natural language tasks. A: Example of dataset SST-ICL task. The example label $y$ is obtained from $v$ by map function $m$, i.e. $y=m(v)$. B: Results on SST-ICL task. We explore the $D_{\text{fix}\to\text{rnd}}\Rightarrow D_{\text{rnd}}$,$D_{\text{fix}\land\text{rnd}}\Rightarrow D_{\text{rnd}}$ and $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$ settings. The dash line denote the time when we transit from $D_{\text{fix}}$ to $D_{\text{rnd}}$. C: Results on WordSelection Task. Adding extra weights loss has significant effect in shortening the learning plateaus.
The previous methods require another data setting $D_{\text{fix}}$ to improve the weights component. However, in some situations, it is difficulty to find such data setting. Therefore, we consider another alternate method by providing extra supervision signal to improve the weights component.
We consider two extra loss setting. The $f^{\prime}_{{\mathbf{w}}^{\prime},s_{c}}(x)$ is denote as the subnetwork of $f_{{\mathbf{w}},s_{c}}(x)$ without the output classifier. We denote the $cls_{e}$ as the classifier for the factor value of factor $e$. The weights loss is defined as
@@
And the context loss is defined as
@@
where $cls_{e_{h}}$ is the classifier to predict hidden factor, and $e_{h}^{(s)}$ is the hidden factor for sequence $s$. Then, the original loss function of Equation 3 is modified into $\mathcal{L}(\theta,s)+\lambda\mathcal{L}_{c}(\theta,s)$ or $\mathcal{L}(\theta,s)+\lambda\mathcal{L}_{w}(\theta,s)$. The $\lambda$ is choosen as $0.1$ in our experiments.
@@ 5.4 Beyond Synthetic Task
SST-ICL task  The task employs the SST dataset as outlined by Socher et al. [41]. A illustration of the task is given in Fig. 6A. Comprehensive details regarding the dataset structure, training methods, and model configurations are available in Appendix B. We assess the results for both the $D_{\text{fix}\to\text{rnd}}\Rightarrow D_{\text{rnd}}$ and $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$ settings. It is noted that improving the weights component through training on $D_{\text{fix}}$ similarly aids in-context learning within this domain. Nevertheless, it should be acknowledged that variations exist between the outcomes of the SST-ICL task and those of the synthetic task; specifically, the enhancements do not surmount the initial learning plateaus. This divergence is attributed to the $D_{\text{fix}}$ condition not being as efficacious in this task as it is in the synthetic one, although it still offers advantages. For example, in the $D_{\text{fix}\to\text{rnd}}$ configuration, adopting a warm-up approach contributes to shortening the second learning plateau. When considering the $D_{\text{rnd}\land\text{fix}}$ setting, a single transition process achieves better performance than the $D_{\text{fix}\to\text{rnd}}$ approach with multiple transitions. The benefits of mix training can be interpreted as the removal of the second or third learning plateaus.
@@
WordSelection task  We have devised an additional task named the WordSelection Task, which requires selecting a single word from a group of four options. For example, given the input “hello information learning art → learning”, the task for the model is to identify “learning” as the correct choice from the provided set. The model must infer the correct answer by considering context examples. We offer five such in-context examples, all selecting the words at the same position as the prompt example. Creating a specific $D_{\text{fix}}$ setting to improve the weights component poses a challenge in this task. Consequently, we opt for the implementation of the extra weights loss technique. We define the factor as the position and the factor valus of a factor is the corresponding word in this position. For example, with ${\mathbf{x}}$ being “hello information learning art”, we define $v_{{\mathbf{x}}}^{(e_{1})}$ as “hello”, $v_{{\mathbf{x}}}^{(e_{2})}$ as “information”, $v_{{\mathbf{x}}}^{(e_{3})}$ as “learning” and $v_{{\mathbf{x}}}^{(e_{4})}$ as “art”. The extra weights loss is then applied in a manner consistent with what is described in Eq. (5). Our experimental findings align with those from the synthetic task, indicating that the advantage of integrating an extra weights loss is not limited to synthetic environments. The results highlight that adding extra weights loss can be a beneficial strategy across different task types.

### v3

@@ 4.3 Dysfunction of weights component
Confusing pattern of weights component  We executed the task under two conditions: $D_{\text{fix}}\Rightarrow D_{\text{fix}}$ and $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$, with the outcomes presented in Fig. 4A. As expected, in the simpler scenario of $D_{\text{fix}}\Rightarrow D_{\text{fix}}$, both the context and weights components exhibit improvement throughout the learning process, leading to enhanced in-context learning performance. However, in the more challenging setting of $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$, the weights component deteriorates over time, with its score remaining below the initial value for the entire duration of the learning process. This differs from the context component, which improves with the rise in in-context learning performance. We refer to the situation where the weights component score falls below the starting value as a dysfunction of the weights component. This outcome is intriguing because training has no effect in improving the weights component. To gain better insight into this phenomenon, we carried out additional experiments across varying levels of task complexity. The model was trained for 50 epochs on these tasks, and we monitored the weights component score post-training. We found that the weights component score gradually declines as the entropy increases, eventually stabilizing at around 0.8, as shown in Fig. 4B.
@@
Weight component degradation is linked to duration learning plateaus.  Our primary concern is the duration of learning plateaus, and we seek to comprehend its connection to the in-weights component. To investigate this, we graphed the relationship between learning plateau length and weights component score after 50 epochs, as shown in Fig. 4C. Our analysis reveals an approximately linear correlation between the weights component score and the learning plateau duration. Short learning plateaus occur when there is a significant improvement in the in-weights component score. Conversely, long learning plateaus arise when the weights component is dysfunctioning or on the edge of dysfunction, that is when the weights component score is at or below the initial value.
@@
Why the weights component is related to learning plateaus.  We are attempting to comprehend this phenomenon, yet theoretically analyzing the learning process of a multilayer transformer on intricate data proves to be a formidable challenge. Recent theoretical studies (Tian et al., 2023; Deora et al., 2023; Huang et al., 2023) focus on the learning dynamics of Transformers with one or two layers using simple datasets. Given these limitations, we propose a more attainable, albeit weaker, construction analysis in Appendix E. Our approach is grounded in the notion that if a model with a good weights component can enhance its in-context learning capabilities with just a few additional parameters, then the weights component must be pivotal for achieving in-context learning prowess. To test this idea, we hypothesize that the in-weights component has been perfectly learned within a specific layer of the Transformer. Our findings reveal that by adding at most three extra Transformer layers specifically tailored for processing contextual information, the model demonstrates significant in-context learning performance. These outcomes suggest that in-context learning abilities are more readily achieved when the weights component is well-optimized.
@@ 5 Breaking Through Learning Plateaus
The previous section examined the mechanisms behind learning plateaus and identified their connection to the weights component. In this section, our goal is to investigate whether we can shorten the learning plateaus or improve the performance increasing for each transition process without scaling models. We have chosen the $D_{\text{rnd}}$ configure as the configuration of the test set by default, as it has demonstrated pronounced learning plateau behavior, and this particular scenario is known to be more effective in evaluating in-context learning capabilities, as discussed by Wei et al. (2023); Min et al. (2022).
@@
Figure 5: Three methods are proposed to assist in overcoming learning plateaus. A: Effective of the warm-up method. Top: Employing $D_{\text{fix}}$ as a warm-up for the Transformer significantly mitigates learning plateaus. The dashed line indicates the transition point from $D_{\text{fix}}$ to $D_{\text{rnd}}$. Bottom: We execute the transition from $D_{\text{fix}\to\text{rnd}}\Rightarrow D_{\text{rnd}}$ at various switching points. The curve labeled ”2” signifies the switch from $D_{\text{fix}}$ to $D_{\text{rnd}}$ at epoch 2. The curve labeled ”0” serves as the baseline, that is, $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$. The dashed lines highlight the respective switching points. B: Combining $D_{\text{fix}}$ and $D_{\text{rnd}}$. Top: Mixed training substantially improves the weights component score during the learning process and eliminates learning plateaus. Bottom: Boosting the weights component can promote the development of in-context learning capabilities in smaller models. The dashed line depicts the task configuration $D_{\text{fix}\land\text{rnd}}\Rightarrow D_{\text{rnd}}$, while the solid line represents the $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$ setting. C: Extra Loss. Top: Incorporating a weights loss can significantly enhance learning, whereas adding context loss does not have a noticeable impact. The baseline is $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$. Bottom: With the weights loss, the Transformer can attain a commendable weights component score after 50 epochs of training. The green dashed line indicates the weights comp. score at the initialization point.
5.1 Weights warm-up method
@@
Previous experiments demonstrate that the model better learns the weight component more effectively in the $D_{\text{fix}}$ setting. A straightforward intuition is whether we can use $D_{\text{fix}}$ to improve the weights component in $D_{\text{rnd}}$ test set. We give a further analysis of the relation between $D_{\text{rnd}}$ and $D_{\text{fix}}$:
@@
Increasing weights component with $D_{\text{fix}}$: Recall that we have ${\mathbb{P}}(y_{p}|{\mathbf{x}}_{p},s_{c})\sim{\mathbb{P}}(v_{p}|{\mathbf{x}}_{p}){\mathbb{P}}(e_{h}|s_{c},m){\mathbb{P}}(m|s_{c})$. Under $D_{\text{fix}}$ setting, we only has one mapping function $m_{0}$. Consequently, our model can readily learn that ${\mathbb{P}}(m_{0}|s_{c})=1$. This simplification allows the model to concentrate on mastering ${\mathbb{P}}(v_{\mathbf{x}}|{\mathbf{x}})$. Therefore, it is anticipated that the model will develop a more refined weights component under this setting. Knowledge transfering between $D_{\text{fix}}$ and $D_{\text{rnd}}$ settings: The knowledge of ${\mathbb{P}}(v_{\mathbf{x}}|{\mathbf{x}})$ is shared between these two settings. The reason is that ${\mathbb{P}}(v_{\mathbf{x}}|{\mathbf{x}})$ is unrelated to the mapping function $m$ and the context $s_{c}$.
@@
Based on this analysis, we propose the following data configuration to improve the weights component on $D_{\text{fix}}$ before training on $D_{\text{rnd}}$. This setting is denoted as $D_{\text{fix}\to\text{rnd}}$, which means that we initially train the model on $D_{\text{fix}}$ for a specific epoch (weights warm-up), and then, we train the model on $D_{\text{rnd}}$. We have the following discovery based on the experiments on $D_{\text{fix}\to\text{rnd}}\Rightarrow D_{\text{rnd}}$.
@@
Weights warm-up helps to mitigate the learning plateaus. From Fig. 5A Top, we observe a notable enhancement in the weights component during the initial “warm-up” phase. Subsequent to this phase, there is a swift improvement in the accuracy of the in-context learning. Nonetheless, a subsequent decline in the weights component score post-warm-up is apparent. This further verify that the Transformer can not learn a good weights component under $D_{\text{rnd}}$ settings. Further investigation is conducted by training the Transformer with various “warm-up” durations under the $D_{\text{fix}}$ setting. As shown in Figure5A Bottom, training the Transformer for a brief number of epochs effectively eliminates the initial learning plateaus. Furthermore, this “warm-up” phase contributes to a reduction in the duration of the subsequent learning plateaus. The application of the “warm-up” technique yields significantly better results compared to those without it, particularly after a training duration of 50 epochs.
5.2 mixed training method
@@
In this section, we try to improve the weights component during the whole training process by designing the $D_{\text{fix}\land\text{rnd}}\Rightarrow D_{\text{rnd}}$ setting. In $D_{\text{fix}\land\text{rnd}}$ setting, half training data comes from $D_{\text{fix}}$ setting, and half training data comes from $D_{\text{rnd}}$ setting. For a fair comparison, the total number of training data is the same as the training set that uses $D_{\text{rnd}}$ or $D_{\text{fix}}$ setting only.
@@
The mixed training method significantly boosts the learning process.  As anticipated, both the weights component and the task accuracy exhibit improvements throughout the learning process, as depicted in Fig.5B Top. And the results do not show a pronounced learning plateau. This suggests that concurrently enhancing the weights component is beneficial. In Fig.5B Bottom, we compare models of various sizes trained on the $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$ setting (represented by the solid line) against those trained on the $D_{\text{rnd}\land\text{fix}}\Rightarrow D_{\text{rnd}}$ setting (depicted with a dashed line). It is observed that a model of size 6.69M trained on $D_{\text{fix}\land\text{rnd}}$, after 20 epochs of training, can attain an accuracy comparable to a model of size 13.88M trained on $D_{\text{rnd}}$ that has undergone 50 epochs of training. These findings suggest that overcoming learning plateaus can lead to a reduction in the computational resources required for training.
@@ 5.3 Extra loss method
The previous methods require another data setting $D_{\text{fix}}$ to improve the weights component. Here, we consider another alternate method by providing an extra supervision signal to improve the weights component when we cannot find the $D_{\text{fix}}$ setting. We consider two extra loss settings. The $f^{\prime}_{{\mathbf{w}}^{\prime},s_{c}}(x)$ is denote as the subnetwork of $f_{{\mathbf{w}},s_{c}}(x)$ without the output classifier. We denote the $cls_{e}$ as the classifier for the factor value of factor $e$. The weights loss is defined as
@@
The context loss is defined as
@@
where $cls_{e_{h}}$ is the classifier to predict hidden factor, and $e_{h}^{(s)}$ is the hidden factor for sequence $s$. Then, the original loss function of Equation 3 is modified into $\mathcal{L}(\theta,s)+\lambda\mathcal{L}_{c}(\theta,s)$ or $\mathcal{L}(\theta,s)+\lambda\mathcal{L}_{w}(\theta,s)$. The $\lambda$ is chosen as $0.1$ in our experiments.
@@
5.4 Further Exploration: Simple Functions Tasks
In this section, we give further exploration of the relation between the weights component and the learning plateaus on the simple function tasks (Garg et al., 2022) and further understand the role of weights component in learning plateaus.
Two Roles of Weights Component  Initially, I’d like to emphasize that the weights component serves two primary functions: 1) It takes the examples $x_{i}$ into the internal working of the Transformer. 2) It aims to find a better representation for $x_{i}$. Considering this, the dysfunction of the weights component can stem from these two sources correspondingly. In the previous analysis of the Shape3D task, the weights component draws from two sources. Since simple function tasks lack representation learning, the dysfunction of the weights component might primarily be attributed to the first reason.
Date generation process  In this task, we generate a dataset with $n$ sequences. The generation process for each sequence is:
  • Initially, we sample a vector ${\mathbf{w}}$ from the Gaussian distribution $\mathcal{N}(0,\mathbf{I}_{d})$ in $\mathbb{R}^{d}$. Each component of ${\mathbf{w}}$ is drawn independently.
  • Subsequently, we sample $\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},……,{\mathbf{x}}_{L}\}$, where each ${\mathbf{x}}_{i}\in\mathbb{R}^{d}$ is drawn from a Gaussian distribution $\mathcal{N}(0,\bm{I}_{d})$.
  • The labels are then determined using the formula $y_{i}=sign({\mathbf{w}}^{T}{\mathbf{x}}_{i})$. Finally, we obtain a sequence $s=\{({\mathbf{x}}_{1},y_{1}),\cdots,({\mathbf{x}}_{L},y_{L})\}$.
Exploring Framework  The training methodology aligns with that described in the paper (Garg et al., 2022). We’ll modify the dimension of ${\mathbf{x}}_{i}$, i.e. the value of $d$ to modulate the difficulty of the task. The evaluation of the weights component follows the same process outlined in Appendix A.3, but with the utilization of the metric $MSE_{p}=\mathbb{E}_{{\mathbf{x}}_{i}}{\|\tilde{{\mathbf{x}}}_{i}-{\mathbf{x}}_{i}\|}$ to assess its performance, where $\tilde{{\mathbf{x}}}$ is the prediction of the probe model. We use the Transformer with 6 layers and we probe at the layer 3. It’s worth noting that unlike the weight component score, where higher is better, lower values of $MSE_{p}$ indicate better weights component.
Figure 6: Experiments results on simple function tasks after 50 epochs. A: the dysfunction of the weights component happens when $d>30$. B: the effect of Extra Loss technique is significant only when the dysfunction of weights component happens.
Results  The $MSE_{p}$ and test accuracy are given in Figure 6. We find that: 1) Dysfunction in the weights component is evident in the SimpleFunction dataset when $d>30$. The Transformer exhibits poor performance in these situations. 2) the effect of the Extra Loss technique is significant only when the dysfunction of weights component happens. The effect is incremental when the Transformer doesn’t experience a significant weights dysfunction. This further verifies the causal relation between the weights component and learning plateaus.
@@ 5.5 Beyond Synthetic Task
Figure 7: Experiments on natural language tasks. A: Example of dataset SST-ICL task. The example label $y$ is obtained from $v$ by map function $m$, i.e. $y=m(v)$. B: Results on SST-ICL task. We explore the $D_{\text{fix}\land\text{rnd}}\Rightarrow D_{\text{rnd}}$ and $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$ settings. The dashed line denotes the time when we transit from $D_{\text{fix}}$ to $D_{\text{rnd}}$. C: Results on WordSelection Task. Adding extra weights loss has a significant effect in shortening the learning plateaus.
@@
SST-ICL task  The task employs the SST dataset as outlined by Socher et al. (2013). An illustration of the task is given in Fig. 7A. Comprehensive details regarding the dataset structure, training methods, and model configurations are available in Appendix B. We assess the results for both the $D_{\text{fix}\land\text{rnd}}\Rightarrow D_{\text{rnd}}$. It is noted that improving the weights component through training on $D_{\text{fix}}$ similarly aids in-context learning within this domain. Nevertheless, it should be acknowledged that variations exist between the outcomes of the SST-ICL task and those of the synthetic task; specifically, $D_{\text{fix}}$ settings worse the performance when trained on small epochs, which is evident in the Figure 7 that the $D_{\text{fix}\land\text{rnd}}$ performs worse than $D_{\text{rnd}}$ when training epoch is less than 30. The key reason to this divergence is that the weights component is more difficult to improve compared that in the synthetic one. As a result, more epochs of training are needed to improve the weights component so as to make it come into effect. This explanation is evident by the result in Figure 7 that the model trained on $D_{\text{fix}\land\text{fix}}$ setting significantly outperforms that trained on $D_{\text{rnd}}$ settings after the 35 epochs.
@@
WordSelection task  We have devised an additional task named the WordSelection Task, which requires selecting a single word from a group of four options. For example, given the input “hello information learning art → learning”, the task for the model is to identify “learning” as the correct choice from the provided set. The model must infer the correct answer by considering context examples. We offer five such in-context examples, all selecting the words at the same position as the prompt example. Creating a specific $D_{\text{fix}}$ setting to improve the weights component poses a challenge in this task. Consequently, we opt for the implementation of the extra weights loss technique. We define the factor as the position and the factor values of a factor is the corresponding word in this position. For example, with ${\mathbf{x}}$ being “hello information learning art”, we define $v_{{\mathbf{x}}}^{(e_{1})}$ as “hello”, $v_{{\mathbf{x}}}^{(e_{2})}$ as “information”, $v_{{\mathbf{x}}}^{(e_{3})}$ as “learning” and $v_{{\mathbf{x}}}^{(e_{4})}$ as “art”. The extra weights loss is then applied in a manner consistent with what is described in Eq. (5). Our experimental findings align with those from the synthetic task, indicating that the advantage of integrating an extra weights loss is not limited to synthetic environments. The results highlight that adding extra weights loss can be a beneficial strategy across different task types.
6 Discussion
Q1: What are the reasons and intuitions behind the dysfunction of weights?  The observed dysfunction in the Transformer’s performance may stem from the confluence of two crucial factors: 1) The problem-solving process relies on the information derived from both context examples and query examples. 2) The contextual information is hard to obtain. In situation 1, where both contextual and weight information are critical, the Transformer is compelled to harness both effectively. Yet, under condition 2, where the contextual information proves difficult to learn, the Transformer finds itself unable to fully extract this vital knowledge. Consequently, it lacks the incentive to refine the weight component at this stage, as doing so would not yield appreciable improvements in overall performance. This phenomenon aligns with our observation in Figure 4C that the weight dysfunction occurs predominantly when the in-context tasks exhibit a high degree of complexity, as measured by the high entropy of the distribution of $m$. As a result, we witness the manifestation of weight dysfunction, characterized by the stagnation or regression of the weight component during the learning process, despite the Transformer’s ongoing attempts to adapt and solve the given problem.
Q2: Whether the decomposition of weights and context components is general?  The conceptual decomposition is general based on the following reasons:
1) Firstly, the decomposition of the weights component and the context component is conceptual instead of physical and only for analysis purposes. The decomposition stems from the understanding that the in-context learning task necessitates information from both its context samples and the query example.
2) Secondly, our evaluation of the weights component and context component is not based on the physical decomposition of these components. We employ complex probe methods (see Section A.3) to analyze these two components because we have only conceptually decomposed them.
3) We tested our method across various scenarios, including the Shape3D task, SST-ICL task, Word Selection Task, and SimpleFunction task, which consistently validated the efficacy of our approach.